Calculate the first two modal frequencies for a fl
Answer #1

here,

For a tube open at both ends,

lamdan = 2L / n

→ where n is the harmonic number,

fundamental n = 1

so

fundamental lamda = 2L = 2 * 0.985 m = 1.97 m

for n = 2

lamda2 = 2 * L/2 = 0.985 m

for speed of sound 340 m/s,

the frequency , f = v/lamda = 172.6 Hz

the frequency 2 , f2 = v/lamda2 = 345.2 Hz